# Supervised Graph Contrastive Learning for Gene Regulatory Networks

### v4

@@ E.4.1 Graph-Level Task
Subtype Classification:
For the subtype classification task, we created a classification model using a 5-class softmax function and trained it using multi-class cross-entropy. Furthermore, performance was evaluated using Accuracy and Macro F1-score. The F1-score results are shown in Appendix F.
E.4.2 Node‐Level Task
BP./CC. Classification :
In Biological Process (BP) classification, three categories for each gene—”metabolism,” ”signal transduction,” and ”cellular organization”—are predicted as a multi-hot vector. In Cellular Component (CC) classification, four categories—”nucleus,” ”mitochondria,” ”endoplasmic reticulum,” and ”plasma membrane”—are predicted as a multi-hot vector. The model was structured using a sigmoid function for each category, and training was performed using binary cross-entropy for each respective category. Performance was evaluated using Subset Accuracy, Macro F1-score, and Jaccard Index as evaluation metrics. The Macro F1-score and Jaccard Index results are shown in Appendix F.
Cancer Rel. Classification
In cancer-related gene classification, 106 genes defined as positive by OncoKB were labeled as ”positive,” and all other genes were labeled as ”negative” for binary classification. A model was created to estimate negative and positive cases using a sigmoid function, and training was performed using binary cross-entropy. Performance was evaluated using Accuracy and F1-score as evaluation metrics. The F1-score results are shown in Appendix F.
Appendix F Additional Results
@@
    Comparison with Augmentation-Adjustment Methods and Augmentation-Free.
@@
    Evaluation on Link Prediction.
Furthermore, an analysis of the robustness of SupGCL against dataset changes is provided in Appendix H.
F.1 Additional Evaluation Metrics
@@
Tables 12 and 13 show the Macro F1-score and Jaccard index results for node-level tasks. Please note that for the cancer-related classification task, we evaluate only the F1-score because it involves binary data. Additionally, Table 14 shows the Macro F1-score for subtype classification in breast cancer. While our proposed method, SupGCL, did not individually achieve state-of-the-art results across all tasks and metrics, it demonstrated the most balanced performance overall.
@@
F.2 Additional Latent Space Analysis
@@
Previously, in Result 3: Latent Space Analysis, we visualized the graph-level embedding space generated by pre-trained models(Figure 3). Node-level results for the breast, lung, and colorectal cancer datasets are presented in Figure 4. Since subtype data were unavailable for the lung and colorectal cancer datasets, only node-level latent space visualizations are presented for these cancers.
These results confirm that both GRACE and our proposed method, SupGCL, yield stable latent representations for these cancers, with no observed latent space collapse. A more detailed analysis on latent space collapse is provided in the next section. Furthermore, a more detailed analysis of the latent space from a biological perspective is provided in Appendix I
@@
To further investigate the characteristics of the node-level latent spaces presented in Result 3: Latent Space Analysis, we employed Principal Component Analysis (PCA). Figure 5 displays the PCA-projected latent spaces from pre-trained models on the breast cancer dataset, along with their corresponding explained variance ratios. For GraphCL, which previously exhibited tendencies towards latent space collapse, this analysis confirmed that its PCA explained variance ratio was overwhelmingly concentrated in the first principal component (PC1), accounting for 98.3%.
@@
F.3 Performance Evaluation across Different Embedding Dimensions
@@
F.4 Statistical Significance Testing
@@
F.5 Comparison with Augmentation-Adjustment Methods and Augmentation-Free
In graph contrastive learning (GCL), topological changes introduced by augmentations such as node or edge dropping are known to degrade performance. To address this issue, recent works have proposed frameworks that either correct or avoid the effects of augmentation. For instance, SGRL (He et al., 2024) (NeurIPS 2024), the successor to BGRL (Thakoor et al. (2021)), has established a state-of-the-art augmentation-free paradigm. Other representative methods include GCA (Zhu et al. (2021)), which automatically adjusts and corrects augmentation effects; AFGRL (Lee et al., 2022), which integrates reconstruction and bootstrap learning; and SimGRACE (Xia et al., 2022), which perturbs model parameters with simple Gaussian noise.
In contrast, SupGCL differs fundamentally from augmentation-free approaches. Rather than avoiding topological changes, SupGCL leverages them as informative signals by supervising augmentations with knockdown-derived GRNs. That is, while conventional approaches aim to circumvent augmentation-induced topology shifts, SupGCL explicitly exploits them as positive supervision, transforming a traditionally negative factor into a beneficial learning signal.
Table 16 reports the comparison on breast cancer GRNs for hazard prediction (graph-level) and BP classification (node-level).
As shown in the results, SupGCL consistently outperformed representative augmentation-correction and augmentation-free models (GCA, AFGRL, SimGRACE). This indicates that actively utilizing topology changes as supervisory signals leads to improved representation learning performance.
Table 16: Comparison with augmentation-correction and augmentation-free methods (Breast Cancer GRN).
Method Hazard (c-index) BP (Subset Accuracy)
GCA $0.620\pm 0.039$ $0.241\pm 0.021$
AFGRL $0.616\pm 0.037$ $0.240\pm 0.026$
SimGRACE $0.638\pm 0.032$ $0.228\pm 0.044$
SupGCL (Ours) $\mathbf{0.650\pm 0.059}$ $\mathbf{0.243\pm 0.052}$
F.6 Evaluation on Link Prediction

### v5

@@ F.5 Robustness to Sampling
The same applies to node-level Graph Contrastive Learning (GCL). For small-scale graphs like Cora or Citeseer, methods such as GRACE utilize all nodes as negative samples, effectively performing $100\%$ sampling. However, for large-scale graphs like Reddit (232,965 nodes), sampling is performed in mini-batch units (e.g., 256 nodes) following GraphSAGE (Hamilton et al., 2017), restricting the ratio to approximately $0.1\%$ ((Xia et al., 2022b). Comparing these trends, the augmentation-wise sampling ratio in SupGCL (8 batch size / $768\sim 948$ node $0.84\sim 1.04\%$) falls within or exceeds the range of general CL methods, indicating that it is at a practically reasonable level from a contrastive learning perspective.
Calculating the order of the error upper bound $\frac{|\mathcal{V}|-|\mathcal{V_{B}}|}{|\mathcal{V_{B}}||\mathcal{V}|}$ based on Proposition F.2 yields the comparisons in Table 13. As shown, the order of the error upper bound for the proposed SupGCL is comparable to that of SimCLR with a batch size of $8,192$ and is sufficiently lower than that of other methods. In other words, in terms of the error upper bound order, the batch size in the augmentation direction employed in this study can be interpreted as sufficiently robust for practical use.
Table 13: Comparison of Sampling Errors Across Methods
Method
No. of Nodes or Samples:
$|\mathcal{V}|$ ($|\mathcal{K}|$)
Batch Size (Node/Sample-level):
$|\mathcal{V_{B}}|$ or $|\mathcal{K_{B}}|$
Sampling Ratio
Order of Sampling deviation
$\frac{|\mathcal{V}|-|\mathcal{V_{B}}|}{|\mathcal{V_{B}}||\mathcal{V}|}$ or $\frac{|\mathcal{K}|-|\mathcal{K_{B}}|}{|\mathcal{K_{B}}||\mathcal{K}|}\cdot\frac{1}{|\mathcal{V}|}$
SimCLR(Chen et al., 2020) $1,281,167$ sample $4096\sim 8192$ sample $0.31\sim 0.63$% $(1.21\sim 2.43)\times 10^{-4}$
Moco v2(He et al., 2020) $1,281,167$ sample $256$ sample 0.010.01% $3.91\times 10^{-3}$
SupCon(Khosla et al., 2020) $1,281,167$ sample $256\sim 512$ sample $0.01\sim 0.03$% $(1.95\sim 3.91)\times 10^{-3}$
ProGCL(Xia et al., 2022b) $232,965$ node $256$ node 0.100.10% $3.90\times 10^{-3}$
SupGCL (Ours) $975(769\sim 946)$ node 88 node $0.84\sim 1.04$% $(1.26\sim 1.27)\times 10^{-4}$
Additionally, regarding the temperature parameter, the settings for SupGCL remain within the scale of existing CL literature. In SimCLR, a temperature of $\tau\approx 0.5$ is standard, and recent reports suggest that lowering it to $\tau\approx 0.25$ can improve performance (Chen et al., 2020). Similarly, MoCo v2 (He et al., 2020) widely adopts $\tau\approx 0.2$. In contrast, the augmentation-wise temperature $\tau_{\rm a}$ used in SupGCL is set in the range of $0.1$ to $2.0$, which is of the same order as existing methods. Therefore, from the perspective of the temperature parameter, the sampling error of SupGCL is expected to exhibit behavior similar to that of conventional CL.
While the practicality of the error upper bound was theoretically confirmed, we further analyzed the impact of increasing the augmentation-wise batch size in additional experiments. Table 14 compares the performance on Hazard and BP prediction tasks for breast cancer patients. We observed an improvement in accuracy as the batch size (more precisely, the sample size for the normalization terms $\hat{C}^{*},\hat{D}^{*}$) increased. However, due to computational resource constraints, we were unable to perform this calculation across all tasks in this study.
Table 14: Downstream task performance vs. Augmentation Batch size
Hazard (Breast) BP (Breast)
$|\mathcal{K_{B}}|=8$ (Setting in paper) 0.650 $\pm$ 0.059 0.243 $\pm$ 0.052
$|\mathcal{K_{B}}|=16$ 0.654 $\pm$ 0.065 0.243 $\pm$ 0.031
$|\mathcal{K_{B}}|=32$ 0.664 $\pm$ 0.052 0.241 $\pm$ 0.043
$|\mathcal{K_{B}}|=40$ 0.663 $\pm$ 0.048 0.251 $\pm$ 0.039
F.6 Proof of Bias Order
The two propositions regarding the error upper bounds presented above are derived from the following lemma on the approximate convergence of KL divergence.
Lemma F.3.
Let the conditional probability distributions $p(j|i)$ and $q(j|i)$ be defined using the Softmax function over a set $V$ as follows:
$p(j|i)$ $=\frac{e^{\beta a_{ij}}}{A_{i}},\quad A_{i}=\sum_{k\in V}e^{\beta a_{ik}},$ (25)
$q(j|i)$ $=\frac{e^{\beta b_{ij}}}{B_{i}},\quad B_{i}=\sum_{k\in V}e^{\beta b_{ik}}.$ (26)
Define the loss function $L$ as the average Kullback-Leibler (KL) divergence:
$L=\frac{1}{|V|}\sum_{i\in V}D_{\mathrm{KL}}(p(\cdot|i)\,\|\,q(\cdot|i)).$ (27)
If $\beta\ll 1$, the loss function can be approximated as:
$L\approx\frac{\beta^{2}}{2|V|}\sum_{i\in V}\mathrm{Var}_{j}(a_{ij}-b_{ij}),$ (28)
where $\mathrm{Var}_{j}(\cdot)$ denotes the variance with respect to the uniform distribution over $j\in V$.
Proof.
We analyze the KL divergence term for a fixed index $i$, denoted as $D_{i}=D_{\mathrm{KL}}(p(\cdot|i)\,\|\,q(\cdot|i))$. By definition:
$D_{i}$ $=\sum_{j\in V}p(j|i)\ln\frac{p(j|i)}{q(j|i)}$
$=\sum_{j\in V}p(j|i)\left((\beta a_{ij}-\ln A_{i})-(\beta b_{ij}-\ln B_{i})\right)$
$=\beta\sum_{j\in V}p(j|i)(a_{ij}-b_{ij})-(\ln A_{i}-\ln B_{i}).$ (29)
Since $\beta\ll 1$, we employ the Taylor expansion for the exponential terms and the logarithm of the partition function. Let $N=|V|$. The partition function $A_{i}$ is expanded as:
$A_{i}$ $=\sum_{j\in V}\left(1+\beta a_{ij}+\frac{\beta^{2}}{2}a_{ij}^{2}+\mathcal{O}(\beta^{3})\right)$
$=N\left(1+\beta\mu_{a_{i}}+\frac{\beta^{2}}{2}m_{a_{i}}^{(2)}+\mathcal{O}(\beta^{3})\right),$ (30)
where $\mu_{a_{i}}=\frac{1}{N}\sum_{j}a_{ij}$ and $m_{a_{i}}^{(2)}=\frac{1}{N}\sum_{j}a_{ij}^{2}$ are the mean and second moment with respect to the uniform distribution. Using the approximation $\ln(1+x)\approx x-\frac{x^{2}}{2}$, the log-partition function becomes:
$\ln A_{i}$ $\approx\ln N+\left(\beta\mu_{a_{i}}+\frac{\beta^{2}}{2}m_{a_{i}}^{(2)}\right)-\frac{1}{2}(\beta\mu_{a_{i}})^{2}$
$=\ln N+\beta\mu_{a_{i}}+\frac{\beta^{2}}{2}\underbrace{(m_{a_{i}}^{(2)}-\mu_{a_{i}}^{2})}_{\sigma_{a_{i}}^{2}},$ (31)
where $\sigma_{a_{i}}^{2}=\mathrm{Var}_{j}(a_{ij})$ under the uniform distribution. Similarly, for $B_{i}$, we have:
$\ln B_{i}\approx\ln N+\beta\mu_{b_{i}}+\frac{\beta^{2}}{2}\sigma_{b_{i}}^{2}.$ (32)
Next, we approximate the expectation term $\sum_{j}p(j|i)(a_{ij}-b_{ij})$. First, approximate $p(j|i)$ up to $\mathcal{O}(\beta)$:
$p(j|i)$ $=\frac{e^{\beta a_{ij}}}{A_{i}}\approx\frac{1+\beta a_{ij}}{N(1+\beta\mu_{a_{i}})}\approx\frac{1}{N}(1+\beta a_{ij})(1-\beta\mu_{a_{i}})$
$\approx\frac{1}{N}\left(1+\beta(a_{ij}-\mu_{a_{i}})\right).$ (33)
Let $\Delta_{ij}=a_{ij}-b_{ij}$. The expectation is:
$\sum_{j\in V}p(j|i)\Delta_{ij}$ $\approx\frac{1}{N}\sum_{j\in V}(1+\beta(a_{ij}-\mu_{a_{i}}))\Delta_{ij}$
$=\frac{1}{N}\sum_{j\in V}\Delta_{ij}+\beta\left(\frac{1}{N}\sum_{j\in V}a_{ij}\Delta_{ij}-\mu_{a_{i}}\frac{1}{N}\sum_{j\in V}\Delta_{ij}\right)$
$=(\mu_{a_{i}}-\mu_{b_{i}})+\beta\mathrm{Cov}_{j}(a_{ij},\Delta_{ij}),$ (34)
where $\mathrm{Cov}_{j}$ denotes covariance under the uniform distribution. Note that $\mathrm{Cov}_{j}(a_{ij},a_{ij}-b_{ij})=\sigma_{a_{i}}^{2}-\mathrm{Cov}_{j}(a_{ij},b_{ij})$.
Substituting equation 31, equation 32, and equation 34 back into equation 29:
$D_{i}$ $\approx\beta\left[(\mu_{a_{i}}-\mu_{b_{i}})+\beta(\sigma_{a_{i}}^{2}-\mathrm{Cov}_{j}(a_{ij},b_{ij}))\right]$
$\quad-\left[(\ln N+\beta\mu_{a_{i}}+\frac{\beta^{2}}{2}\sigma_{a_{i}}^{2})-(\ln N+\beta\mu_{b_{i}}+\frac{\beta^{2}}{2}\sigma_{b_{i}}^{2})\right].$ (35)
The first-order terms $\beta(\mu_{a_{i}}-\mu_{b_{i}})$ cancel out. Collecting the $\beta^{2}$ terms:
$D_{i}$ $\approx\beta^{2}\left(\sigma_{a_{i}}^{2}-\mathrm{Cov}_{j}(a_{ij},b_{ij})-\frac{1}{2}\sigma_{a_{i}}^{2}+\frac{1}{2}\sigma_{b_{i}}^{2}\right)$
$=\frac{\beta^{2}}{2}\left(\sigma_{a_{i}}^{2}-2\mathrm{Cov}_{j}(a_{ij},b_{ij})+\sigma_{b_{i}}^{2}\right)$
$=\frac{\beta^{2}}{2}\mathrm{Var}_{j}(a_{ij}-b_{ij}).$ (36)
Finally, averaging over all $i\in V$:
$L=\frac{1}{|V|}\sum_{i\in V}D_{i}\approx\frac{\beta^{2}}{2|V|}\sum_{i\in V}\mathrm{Var}_{j}(a_{ij}-b_{ij}).$ (37)
Lemma F.4.
Define the KL divergence within a batch $U\subset V$ as:
$\hat{p}(j|i)$ $=\frac{1}{\hat{A}_{i}}e^{\beta a_{ij}},\quad\hat{A}_{i}=\sum_{j\in U}e^{\beta a_{ij}}$
$\hat{q}(j|i)$ $=\frac{1}{\hat{B}_{i}}e^{\beta b_{ij}},\quad\hat{B}_{i}=\sum_{j\in U}e^{\beta b_{ij}}$
$\hat{L}$ $=\frac{1}{|U|}\sum_{i\in U}D_{\mathrm{KL}}(\hat{p}(j|i)|\hat{q}(j|i))$
Then, the following approximation holds:
$L-\mathbb{E}_{B}[\hat{L}]\approx\frac{|V|-|U|}{(|V|-1)|U|}\cdot\frac{\beta^{2}}{2}\cdot\frac{1}{|V|}\sum_{i\in V}\mathrm{Var}_{j}(a_{ij}-b_{ij}).$ (38)
Proof.
Let $N=|V|$ and $M=|U|$. By applying the result of the previous Lemma to the batch $U$, the batch loss $\hat{L}$ for a specific batch $U$ can be approximated as:
$\hat{L}\approx\frac{\beta^{2}}{2M}\sum_{i\in U}\mathrm{Var}_{j\in U}(\Delta_{ij}),$ (39)
where $\Delta_{ij}=a_{ij}-b_{ij}$, and $\mathrm{Var}_{j\in U}(\cdot)$ denotes the variance over the set $U$ (sample variance). We consider the expectation of $\hat{L}$ with respect to the random choice of the batch $U$ (sampled uniformly without replacement from $V$). By the linearity of expectation:
$\mathbb{E}_{U}[\hat{L}]\approx\frac{\beta^{2}}{2}\mathbb{E}_{U}\left[\frac{1}{M}\sum_{i\in U}\hat{\sigma}_{i}^{2}(U)\right],$ (40)
where $\hat{\sigma}_{i}^{2}(U)=\mathrm{Var}_{j\in U}(\Delta_{ij})$. We can rewrite the expectation using an indicator variable $\mathbb{I}(i\in U)$:
$\mathbb{E}_{U}\left[\frac{1}{M}\sum_{i\in V}\mathbb{I}(i\in U)\hat{\sigma}_{i}^{2}(U)\right]$ $=\frac{1}{M}\sum_{i\in V}P(i\in U)\mathbb{E}[\hat{\sigma}_{i}^{2}(U)\mid i\in U]$
$=\frac{1}{M}\sum_{i\in V}\frac{M}{N}\mathbb{E}[\hat{\sigma}_{i}^{2}(U)\mid i\in U]$
$=\frac{1}{N}\sum_{i\in V}\mathbb{E}[\hat{\sigma}_{i}^{2}(U)\mid i\in U].$
Here, $\mathbb{E}[\hat{\sigma}_{i}^{2}(U)\mid i\in U]$ represents the expected variance of the values ${\Delta_{ij}},~{j\in U}$ when the batch $U$ is a random subset of size $M$ containing $i$. Since the variance is computed over $j\in U$, and $U$ is drawn from $V$, we invoke the relationship between the expected sample variance (biased estimator) and the population variance. Let $\sigma_{i}^{2}=\mathrm{Var}_{j\in V}(\Delta_{ij})$ be the population variance. The expectation of the sample variance defined with denominator $M$ is given by:
$\mathbb{E}[\hat{\sigma}_{i}^{2}(U)]=\frac{M-1}{M}\frac{N}{N-1}\sigma_{i}^{2}.$ (41)
Substituting this back into the expression for $\mathbb{E}_{U}[\hat{L}$]:
$\mathbb{E}_{U}[\hat{L}$] $\approx\frac{\beta^{2}}{2N}\sum_{i\in V}\left(\frac{M-1}{M}\frac{N}{N-1}\sigma_{i}^{2}\right)$
$=\frac{M-1}{M}\frac{N}{N-1}\left(\frac{\beta^{2}}{2N}\sum_{i\in V}\sigma_{i}^{2}\right)$
Finally, we compute the difference $L-\mathbb{E}_{U}[\hat{L}$]:
$L-\mathbb{E}_{U}[\hat{L}$] $\approx\left(1-\frac{N(M-1)}{M(N-1)}\right)\left(\frac{\beta^{2}}{2N}\sum_{i\in V}\sigma_{i}^{2}\right)$
$=\frac{N-M}{M(N-1)}\left(\frac{\beta^{2}}{2N}\sum_{i\in V}\sigma_{i}^{2}\right)$
This completes the proof. ∎
The proof above similarly holds for the case where $p(j|i)=\delta_{ij}$, resulting in an error order of $O(\beta^{2}\cdot\tfrac{|V|-|U|}{|V||U|})$. This result supports Proposition F.2 regarding the error order due to node sampling.
Crucially, for the error upper bound of SupGCL, the term $\mathrm{Var}_{j}(a_{ij}-b_{ij})$ plays a significant role. In the node-level probability density function, $a_{ij},b_{ij}$ correspond to the similarity between nodes $\langle z^{*}_{i}|z_{j}^{*}\rangle$, whereas in the augmentation-level probability density function, $a_{ij},b_{ij}$ represent the similarity between graphs $\mathrm{sim}_{F}(Z^{a},Z^{b})$. Assuming that the similarity of each node is sampled from an identical distribution, i.e.,
$\langle z_{i}^{a}|z_{i}^{b}\rangle\sim P^{a,b},$ (42)
the graph-level similarity can be regarded as the sample mean of the node-level similarities. Therefore, the variance of the graph-level similarity is given by:
$\mathrm{Var}(\mathrm{sim}_{F}(Y^{a},Y^{b})-\mathrm{sim}_{F}(Z^{a},Z^{b}))$ $=\mathrm{Var}\left(\frac{1}{|V|}\sum_{i\in V}\langle y_{i}^{a},y_{i}^{b}\rangle-\frac{1}{|V|}\sum_{i\in V}\langle z_{i}^{a},z_{i}^{b}\rangle\right)$
$=\frac{1}{|V|}\mathrm{Var}(\langle y^{a},y^{b}\rangle-\langle z^{a},z^{b}\rangle),$
where $z^{*},y^{*}$ are random variables representing node-level features. Consequently, the order of the error upper bound for SupGCL satisfies $O(\beta^{2}\cdot\tfrac{|V|-|U|}{|V||U|}\cdot\tfrac{1}{|V|})$, confirming the validity of Proposition F.1.
In this proof, we adopt the i.i.d. assumption as shown in Eq. 42. This constitutes a strong assumption that the inner products of node-level features follow a probability distribution that is independent of the nodes. However, by employing Hoeffding’s inequality, we can relax this assumption due to the boundedness of the feature inner products. Consequently, the relaxed assumption requires only the independence of the node-level distributions, which is the condition for Hoeffding’s inequality.
Appendix G Additional Results
@@
    Evaluation on Link Prediction.
@@
    Comparison of Data Augmentation Strategies
Furthermore, an analysis of the robustness of SupGCL against dataset changes is provided in Appendix I.
G.1 Additional Evaluation Metrics
@@
Tables 15 and 16 show the Macro F1-score and Jaccard index results for node-level tasks. Please note that for the cancer-related classification task, we evaluate only the F1-score because it involves binary data. Additionally, Table 17 shows the Macro F1-score for subtype classification in breast cancer. While our proposed method, SupGCL, did not individually achieve state-of-the-art results across all tasks and metrics, it demonstrated the most balanced performance overall.
@@
G.2 Additional Latent Space Analysis
@@
Previously, in Sec 5.2, we visualized the graph-level embedding space generated by pre-trained models (Figure 3). Node-level results for the breast, lung, and colorectal cancer datasets are presented in Figure 4. Since subtype data were unavailable for the lung and colorectal cancer datasets, only node-level latent space visualizations are presented for these cancers.
These results confirm that both GRACE and our proposed method, SupGCL, yield stable latent representations for these cancers, with no observed latent space collapse. A more detailed analysis on latent space collapse is provided in the next section. Furthermore, a more detailed analysis of the latent space from a biological perspective is provided in Appendix J
@@
To further investigate the characteristics of the node-level latent spaces presented in Sec 5.2, we employed Principal Component Analysis (PCA). Figure 5 displays the PCA-projected latent spaces from pre-trained models on the breast cancer dataset, along with their corresponding explained variance ratios. For GraphCL, which previously exhibited tendencies towards latent space collapse, this analysis confirmed that its PCA explained variance ratio was overwhelmingly concentrated in the first principal component (PC1), accounting for 98.3%.
@@
G.3 Performance Evaluation across Different Embedding Dimensions
@@
G.4 Statistical Significance Testing
@@
G.5 Evaluation on Link Prediction